OpenStudy (eric_d):

Find J(x) if x^3+3x^2+x-2=(x+2).J(x) @ikram002p

OpenStudy (eric_d):

the given answer is x^2+x-1

OpenStudy (ikram002p):

x^3+3x^2+x-2=(x+2).J(x) so j(x) =x^3+3x^2+x-2 / (x+2) use long division

OpenStudy (ikram002p):



OpenStudy (eric_d):

Thnx @ikram002p

OpenStudy (ikram002p):

np :)